# Optimised neural networks for online processing of ATLAS calorimeter data on FPGAs

### v2

@@ 1 Introduction
New neural network–based algorithms have been investigated to replace the OF algorithm for LAr data processing [LArNN1]. These algorithms have been shown to outperform the OF in terms of energy resolution in events with overlapping pulses due to pile-up, within an energy range between $0\text{\,}$ and $5\text{\,}\mathrm{GeV}$. For deployment in the LAr calorimeter readout, the neural networks must satisfy stringent requirements on latency and network size to match the computing resources available on the FPGAs, while still achieving a cell energy resolution superior to that of the OF. Neural networks with reduced size, implementing fewer than 500 multiply-accumulate operations (MAC units), have already been successfully implemented on FPGAs [LArNN2].
@@ 2 Description of the network architectures
The training is performed with the Tensorflow Keras toolkit [tensorflow]. One million events are used for training, 1.5 millions for validation, and 2.5 millions for testing. Since the Bayesian optimisation (Section 3) uses the test sample to tune hyperparameters, an independent sample of 13 million events is used for the final evaluation. The detailed architectures and parameters of the networks are summarised in Table 1.
@@ 3.1 Hardware-constrained objective function
No penalty on the network size is applied below 500 MAC units, a linear penalty is introduced above 500, and an additional exponential penalty beyond 1200, where FPGA implementation becomes increasingly difficult. The constant parameters in Equation 1 are chosen to ensure that the function is continuous and are tuned heuristically to achieve the best performance. Since neural network training is not deterministic and depends on random weight initialisation, five trainings are performed for each parameter set, and the configuration with the best objective score is selected.

### v3

@@ 1 Introduction
New neural network–based algorithms have been investigated to replace the OF algorithm for LAr data processing [LArNN1]. These algorithms have been shown to outperform the OF in terms of energy resolution in events with overlapping pulses due to pile-up, within an energy range between $0\text{\,}$ and $5\text{\,}\mathrm{GeV}$. For deployment in the LAr calorimeter readout, the neural networks must satisfy stringent requirements on latency and network size to match the computing resources available on the FPGAs, while still achieving a cell energy resolution superior to that of the OF. Neural networks of reduced size, implementing around 300 multiply–accumulate operations (MAC units), have already been successfully deployed on previous-generation INTEL FPGAs [LArNN2]. Given the increased resources and operating frequency of the Agilex 7 FPGA, a neural network with fewer than 500 MAC units is well within the capabilities of the LASP boards.
@@ 2 Description of the network architectures
The training is performed with the Tensorflow Keras toolkit [tensorflow]. One million events are used for training, 1.5 million for validation, and 2.5 million for testing. The training sample size was found to be sufficient to reach optimal performance. Since the Bayesian optimisation (Section 3) uses the test sample to tune the hyperparameters, an independent sample of 13 million events is used for the final evaluation. The size of this latter sample was chosen to minimise statistical fluctuations across the transverse-energy ranges in which the networks are assessed. The detailed architectures and parameters of the networks are summarised in Table 1.
@@ 3.1 Hardware-constrained objective function
No penalty on the network size is applied below 500 MAC units, a linear penalty is introduced above 500, and an additional exponential penalty beyond 850, where FPGA implementation becomes increasingly difficult. The constant parameters in Equation 1 are chosen to ensure that the function is continuous and are tuned heuristically to achieve the best performance. Since neural network training is not deterministic and depends on random weight initialisation, five trainings are performed for each parameter set, and the configuration with the best objective score is selected.